# A Review of Modern Recommender Systems Using Generative Models (Gen-RecSys)

### v1

@@ Abstract.
Traditional recommender systems (RS) have used user-item rating histories as their primary data source, with collaborative filtering being one of the principal methods. However, generative models have recently developed abilities to model and sample from complex data distributions, including not only user-item interaction histories but also text, images, and videos – unlocking this rich data for novel recommendation tasks. Through this comprehensive and multi-disciplinary survey, we aim to connect the key advancements in RS using Generative Models (Gen-RecSys), encompassing: a foundational overview of interaction-driven generative models; the application of large language models (LLM) for generative recommendation, retrieval, and conversational recommendation; and the integration of multimodal models for processing and generating image and video content in RS. Our holistic perspective allows us to highlight necessary paradigms for evaluating the impact and harm of Gen-RecSys and identify open challenges. A more up-to-date version of the papers are maintained at: https://github.com/yasdel/LLM-RecSys.
Generative Models, Recommender Systems, GANs, VAEs, LLMs, Personalization, Ethical and Societal Considerations
@@ 1. Introduction
The core of generative models lies in their ability to model and sample from the data distribution they are trained on for various inferential purposes. This enables two primary modes of application for RS:
@@
    Directly trained models. This approach trains generative models, such as VAE-CF(Variational AutoEncoders for Collaborative Filtering) (Liang et al., 2018), directly on user-item interaction data to predict user preferences, without using large, diverse pretraining datasets. These models learn the probability distribution of items a user might like based on their previous interactions.
@@
    Pretrained models. This strategy uses models pretrained on diverse data (text, images, videos) to understand complex patterns, relationships, and contexts. Among other methods, this survey covers the use of pretrained generative models for RS in the following settings:
    • Zero- and Few-shot Prompting (cf. Section 3.2.1), with the latter using in-context learning (ICL) for broad understanding without extra training.
    • Fine-Tuning (cf. Section 3.3), adjusting models with specific datasets for tailored recommendations.
@@
    • Embedding for Downstream Training (cf. Section 3.1), generating embeddings or token sequences for complex content representation.
    • Multimodal Approaches (cf. Section 4), using various data types to improve the accuracy and relevance of model recommendation.
@@
Leveraging text in RS, including descriptions of items and user preferences as well as user-system interactions such as reviews, queries or conversation histories, with examples including generative recommendation and explanation, and conversational recommendation. Sections 3.2, 3.5
@@
Selecting or generating informative user-item interactions to improve recommendation model training. Sections 2.3, 2.4
@@ 1.1. Recent Surveys and Our Contributions
  • Lin et al. (2023) focus on adapting LLMs in RS, detailing various tasks and applications.
  • Fan et al. (2023) overview LLMs in RS, emphasizing pre-training, fine-tuning, and prompting.
  • Vats et al. (2024) review LLM-based RS, introducing a heuristic taxonomy for categorization.
@@
While the mentioned surveys offer crucial insights, their scope is often limited to LLMs (Li et al., 2023b; Wu et al., 2023b; Lin et al., 2023; Fan et al., 2023; Vats et al., 2024) or, more broadly, FMs (Huang et al., 2024a) and/or specific sets of models such as GANs (Deldjoo et al., 2021), without encompassing the wider spectrum of generative models and data modalities. In contrast, the work by (Wang et al., 2023b) provides a more comprehensive survey focused on these broader aspects of Gen-RecSys although their work is mostly on personalized content generation.
@@
Figure 1 depicts the organization of our survey on Gen-RecSys. It is categorized into layers that explores data sources, recommendation models, and scenarios with a vertical hierarchy that extends to the evaluation of these systems and the challenges they face. We thus offer a systematic approach to breaking down the recommendation process of Gen-RecSys research into discrete components and methodologies. Our contributions can be briefly summarized as follows.
@@
    We study the evaluation of Gen-RecSys (cf. Section 5) with finer details, shedding light on multiple aspects such as benchmarks, evaluation for impact and harm relative to multiple stakeholders, and conversational evaluation. This evaluation framework is particularly notable as it helps to understand the complex challenges intrinsic to Gen-RecSys.
@@
    We discuss several open research challenges and issues (cf. Section 6). Our survey benefits from the expertise of distinguished scholars and industry practitioners from a variety of institutions and disciplines.
@@ 2. Generative Models for Interaction-
Interaction-driven recommendation is a setup where only the user-item interactions (e.g., “user A clicks item B”) are available, which is the most general setup studied in RS. In this setup, we concentrate on the inputs of user-item interactions and outputs of item-recommended lists or grids rather than richer inputs or outputs from other modalities such as recommendation explanations. Even though no textual or visual information is involved, Deep Generative Models (DGMs) (Goodfellow et al., 2014; Kingma and Welling, 2013; Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017; Sohl-Dickstein et al., 2015) still show their unique usefulness. For example, DGMs can augment user-item interactions for model training (Cai and Wang, 2018), use denoising (Wang et al., 2023d) for recommendation, and learn the distributions of recommendation layouts (Xiao et al., 2022).
In this section, we examine the paradigms of DGMs for recommendation tasks with data on user-item interactions, including auto-encoding models (Kingma and Welling, 2013), auto-regressive models (Vaswani et al., 2017; Hochreiter and Schmidhuber, 1997), generative adversarial networks (Goodfellow et al., 2014), diffusion models (Sohl-Dickstein et al., 2015) and more [^1] When discussing DGMs, confusion may arise due to the conflation of model paradigms and architectures. Model architectures, such as multilayer perceptron (Hornik et al., 1989), self-attentive models (Vaswani et al., 2017) can be used across various paradigms (Kang and McAuley, 2018; Liu et al., 2023a)..
@@ 2.1.1. Preliminaries: Denoising Auto-Encoding Models
Denoising Auto-Encoding models are a group of models that learn to recover the original inputs from a corrupted version of the inputs. Traditionally, denoising auto-encoding models refer to a group of Denoising Autoencoders (Vincent et al., 2008; Sedhain et al., 2015) with hidden layers as a “bottleneck”. For example, AutoRec (Sedhain et al., 2015) tries to reconstruct the input vector, which is partially observed. More broadly, BERT-like models (Devlin et al., 2019a; Sun et al., 2019; Wu et al., 2016) are also treated as denoising auto-encoding models. Such models recover corrupted (i.e., masked) inputs through stacked self-attention blocks (Sun et al., 2019; He et al., 2021). For example, BERT4Rec (Sun et al., 2019) is trained to predict masked items in given user historical interaction sequences. Therefore, BERT-like (Devlin et al., 2019a) models can be used for next-item prediction in the inference phase (Sun et al., 2019; He et al., 2021).
@@ 2.1.2. Variational Auto-Encoding Models
Variational Autoencoders (VAEs) are models that learn stochastic mappings from an input $x$ from a often complicated probability distribution $p$ to a probability distribution $q$. This distribution, $q$, is typically simple (e.g., a normal distribution), enabling the use of a decoder to generate outputs $\hat{x}$ by sampling from $q$ (Kingma and Welling, 2013). VAEs find wide application in traditional RS, particularly for collaborative filtering (Liang et al., 2018), sequential recommendation (Sachdeva et al., 2019) and slate generation (Jiang et al., 2018; Liu et al., 2021; Deffayet et al., 2023). Compared to Denoising Autoencoders (DAEs), VAEs often demonstrate superior performance in collaborative filtering due to stronger modeling assumptions, such as VAE-CF (Liang et al., 2018). Additionally, Conditional VAE (CVAE) (Sohn et al., 2015) models learn distributions of preferred recommendation lists for a given user. This makes them useful for generating those lists beyond a greedy ranking schema. Examples like ListCVAE (Jiang et al., 2018) and PivotCVAE (Liu et al., 2021) use VAEs to generate entire recommendation lists rather than solely ranking individual items.
@@ 2.2.2. Self-Attentive Auto-Regressive Models
Self-attentive models replace the recurrent unit with self-attention and related modules, inspired by transformers (Vaswani et al., 2017). This group of models can be used in session-based recommendation and sequential recommendation (Kang and McAuley, 2018; Wu et al., 2020; Lin et al., 2020), next-basket or bundle prediction (Yu et al., 2020), and model attacking (Yue et al., 2021), similar to RNNs. Meanwhile, the benefits of self-attentive models are that they handle long-term dependencies better than RNNs and enable parallel training (Vaswani et al., 2017). Additionally, self-attentive models are the de-facto option for pre-trained models (Devlin et al., 2019a) and large language models (Brown et al., 2020; Wei et al., 2022; Bubeck et al., 2023), which is gaining traction in RS. More details about using such language models for recommendations will be discussed in Section 3.
@@ 2.4. Diffusion Models
Diffusion models  (Sohl-Dickstein et al., 2015) generate outputs through a two-step process: (1) corrupting inputs into noise via a forward process, and (2) learning to recover the original inputs from the noise iteratively in a reverse process. Their impressive generative capabilities have attracted growing interest from the recommender systems community.
First, a group of works (Wang et al., 2023d; Walker et al., 2022) aims to learn users’ future interaction probabilities through diffusion models. For example, DiffRec (Wang et al., 2023d) predicts users’ future interactions using corrupted noises from the users’ historical interactions. Second, another group of works (Liu et al., 2023d; Wu et al., 2023a) focuses on using diffusion models to augment training sequences, showing promising results in alleviating the data sparsity and long-tail user problems in sequential recommendation.
@@ 3. Large Language Models in Recommendation
While language has been leveraged by content-based RS for over three decades (Lops et al., 2011), the advent of pretrained LLMs and their emergent abilities for generalized, multi-task natural language (NL) reasoning (Brown et al., 2020; Wei et al., 2022; Bubeck et al., 2023) has marked the beginning of a new stage of language-based recommendation. Exploiting textual content is a compelling prospect, since NL constitutes a unified, expressive, and interpretable medium that can represent not only item features or user preferences, but also user-system interactions, recommendation tasks, and external knowledge. Specifically, real items are often associated with rich text content including titles, descriptions, semi-structured textual metadata, and reviews. Similarly, user preferences can be easily articulated in NL in both traditional systems and the emerging paradigm of conversational RS: in the former by leveraging NL components such as reviews, search queries, or liked item descriptions, and in the latter through direct NL user utterances. Pretrained LLMs provide new ways to exploit this textual data, since, in many domains, LLMs have learned useful reasoning abilities for making and explaining item recommendations based on user preferences (Sanner et al., 2023; Sileo et al., 2022; Geng et al., 2022; He et al., 2023). These pretrained abilities can be further augmented through prompting (Sanner et al., 2023; Sileo et al., 2022; Liu et al., 2023c), fine-tuning (Geng et al., 2022; Bao et al., 2023; Harte et al., 2023; Mao et al., 2023; Kang et al., 2023; Cui et al., 2022; Zhang et al., 2023b), retrieval (Hou et al., 2023; Chen, 2023; Wang and Lim, 2023; Dai et al., 2023), and external tools (Wang et al., 2023a; Friedman et al., 2023; Gao et al., 2023).
We next explore key evolving approaches using LLMs in RS in both single- and multi-stage settings, covering:
  • Encoder-only LLM Recommendation: Section 3.1 discusses LLMs used either as highly-scalable dense retrievers or to cross-encode preference and item information for top-K recommendation and rating prediction.
  • LLM-based Generative Recommendation Section 3.2 introduces generative recommendation with LLMs in zero-shot (ZS) and few-shot (FS) settings, as well as approaches to internalize recommendation knowledge through fine-tuning, prompt-tuning, and controllable decoding.
  • LLM/RS Hybrids We explore the complimentary use of LLMs and RS in multi-stage pipelines, with Section 3.3 focusing on retrieval-augmented generation and Section 3.4 discussing LLMs as input generators for RS.
  • Conversational Recommendation: Section 3.5 covers recommendation in NL dialogue, including work on tool augmented LLM systems for of dialogue management, recommendation, explanation, and retrieval.
@@ 3.1.1. Recommendation as Dense Retrieval
To use dense retrieval for recommendation, each item’s text content (e.g., title, description, and/or reviews) is treated as a document, and if no explicit query is available from the user, a common approach is to synthesize a query by concatenating the descriptions of a user’s recently liked items (often titles for conciseness (Harte et al., 2023; Penha and Hauff, 2020; Li et al., 2023c)). Dense retrievers (e.g., BERT (Devlin et al., 2019b), TAS-B (Hofstätter et al., 2021), Condenser (Gao and Callan, 2021)) produce a ranked list of documents given a query by evaluating the similarity (e.g., dot product or cosine similarity) between an encoder-only LLM document embedding and query embedding. Dense retrieval is highly scalable (especially with approximate search libraries like FAISS [^2] https://github.com/facebookresearch/faiss) as documents and queries are encoded separately, allowing for dense vector indexing of documents before querying. Recent work explores recommendation as dense retrieval with off-the-shelf retrievers (Penha and Hauff, 2020; Harte et al., 2023), contrastive BERT fine-tuning for review-based retrieval (Abdollah Pour et al., 2023), the use of a second-level encoder fusing the embedding of a user’s recently liked items into a user embeddings before scoring (Wu et al., 2019; Li et al., 2022b) or fine-tuned components of multi-stage systems (Mysore et al., 2023; Li et al., 2023c; Hou et al., 2023).

### v2

@@ Abstract.
Traditional recommender systems typically use user-item rating histories as their main data source. However, deep generative models now have the capability to model and sample from complex data distributions, including user-item interactions, text, images, and videos, enabling novel recommendation tasks. This comprehensive, multidisciplinary survey connects key advancements in RS using Generative Models (Gen-RecSys), covering: interaction-driven generative models; the use of large language models (LLM) and textual data for natural language recommendation; and the integration of multimodal models for generating and processing images/videos in RS. Our work highlights necessary paradigms for evaluating the impact and harm of Gen-RecSys and identifies open challenges. This survey accompanies a tutorial presented at ACM KDD’24, with supporting materials provided at: https://encr.pw/vDhLq.
Generative Models, Recommender Systems, GANs, VAEs, LLMs, Multimodal, vLLMs, Ethical and Societal Considerations
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671474isbn: 979-8-4007-0490-1/24/08ccs: Information systems Recommender systems
@@ 1. Introduction
The core of generative models lies in their ability to model and sample from their training data distribution for various inferential purposes, which enables two primary modes of application for RS:
@@
    Directly trained models. This approach trains generative models, such as VAE-CF (Variational AutoEncoders for Collaborative Filtering) (Liang et al., 2018) (cf. Section 2.1) directly on user-item interaction data to predict user preferences, without using large, diverse pretraining datasets. These models learn the probability distribution of items a user might like based on their previous interactions.
@@
    Pretrained models. This strategy uses models pretrained on diverse data (text, images, videos) to understand complex patterns, relationships, and contexts that often exhibit (emergent) generalization abilities to a range of novel tasks (Wei et al., 2022). Among a variety of applications, this survey covers the use of pretrained Gen-RecSys models in the following settings:
    • Zero- and Few-shot Learning (cf. Section 3.2.1), using in-context learning (ICL) for broad understanding without extra training.
    • Fine-Tuning (cf. Section 3.3), adjusting model parameters using specific datasets for tailored recommendations.
@@
    • Feature Extraction for Downstream Recommendation (cf. Section 3.4), e.g., generating embeddings or token sequences for complex content representation.
    • Multimodal Approaches (cf. Section 4), jointly using multiple data types such as text, image, and video to enhance and improve the recommendation experience.
@@
Utilize text in RS, including item descriptions, user preferences, reviews, queries, and conversation histories. Examples include generative and conversational recommendations and explanations. Sections 3.2, 3.5
@@
Selecting or generating informative user-item interactions to improve RS model training. Sections 2.3, 2.4
@@ 1.1. Recent Surveys and Our Contributions
  • Lin et al. (2023) focus on adapting LLMs in RS, detailing various tasks and applications. Fan et al. (2023) overview LLMs in RS, emphasizing pre-training, fine-tuning, and prompting, while Vats et al. (2024) review LLM-based RS, introducing a heuristic taxonomy for categorization.
@@
While the mentioned surveys offer crucial insights, their scope is often limited to LLMs (Li et al., 2023b; Wu et al., 2023b; Lin et al., 2023; Fan et al., 2023; Vats et al., 2024) or, more broadly, FMs (Huang et al., 2024a) and/or specific models such as GANs (Deldjoo et al., 2021), without considering the wider spectrum of generative models and data modalities. The work by (Wang et al., 2023b) provides a more relevant survey on Gen-RecSys although their work is mostly on personalized content generation.
@@
Figure 1 illustrates the structure of our Gen-RecSys survey. It categorizes data sources, recommendation models, and scenarios, extending to system evaluation and challenges. We present a systematic approach to deconstructing the Gen-RecSys recommendation process into distinct components and methodologies. Our contributions are summarized as follows.
@@
    We study the evaluation of Gen-RecSys with finer details, shedding light on multiple aspects such as benchmarks, evaluation for impact and harm relative to multiple stakeholders, and conversational evaluation. This evaluation framework is particularly notable as it helps to understand the complex challenges intrinsic to Gen-RecSys.
@@
    We discuss several open research challenges and issues. Our survey benefits from the expertise of scholars/industry practitioners from diverse institutions and disciplines.
@@ 2. Generative Models for Interaction-
Interaction-driven recommendation is a setup where only the user-item interactions (e.g., “user A clicks item B”) are available, which is the most general setup studied in RS. In this setup, we concentrate on the inputs of user-item interactions and outputs of item-recommended lists or grids rather than richer inputs or outputs from other modalities such as textual reviews. Even though no textual or visual information is involved, generative models (Goodfellow et al., 2014; Kingma and Welling, 2013; Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017; Sohl-Dickstein et al., 2015) still show their unique usefulness. In this section, we examine the paradigms of generative models for recommendation tasks with user-item interactions, including auto-encoding models (Kingma and Welling, 2013), auto-regressive models (Vaswani et al., 2017; Hochreiter and Schmidhuber, 1997), generative adversarial networks (Goodfellow et al., 2014), diffusion models (Sohl-Dickstein et al., 2015) and more.
@@ 2.1.1. Preliminaries: Denoising Auto-Encoding Models
Denoising Auto-Encoding models are a group of models that learn to recover the original inputs from a corrupted version of the inputs. Traditionally, denoising auto-encoding models refer to a group of Denoising Autoencoders (Vincent et al., 2008; Sedhain et al., 2015) with hidden layers as a “bottleneck”. For example, AutoRec (Sedhain et al., 2015) tries to reconstruct the input vector, which is partially observed. More broadly, BERT-like models (Devlin et al., 2019; Sun et al., 2019; Wu et al., 2016) are also treated as denoising auto-encoding models. Such models recover corrupted (i.e., masked) inputs through stacked self-attention blocks (Sun et al., 2019; He et al., 2021). For example, BERT4Rec (Sun et al., 2019) is trained to predict masked items in given user historical interaction sequences. Therefore, BERT-like (Devlin et al., 2019) models can be used for next-item prediction in the inference phase (Sun et al., 2019; He et al., 2021).
@@ 2.1.2. Variational Auto-Encoding Models
Variational Autoencoders (VAEs) are models that learn stochastic mappings from an input $x$ from a often complicated probability distribution $p$ to a probability distribution $q$. This distribution, $q$, is typically simple (e.g., a normal distribution), enabling the use of a decoder to generate outputs $\hat{x}$ by sampling from $q$ (Kingma and Welling, 2013). VAEs find wide applications in traditional RS, particularly for collaborative filtering (Liang et al., 2018), sequential recommendation (Sachdeva et al., 2019) and slate generation (Jiang et al., 2018; Liu et al., 2021; Deffayet et al., 2023). Compared to Denoising Autoencoders, VAEs often demonstrate superior performance in collaborative filtering due to stronger modeling assumptions, such as VAE-CF (Liang et al., 2018). Additionally, Conditional VAE (CVAE) (Sohn et al., 2015) models learn distributions of preferred recommendation lists for a given user. This makes them useful for generating those lists beyond a greedy ranking schema. Examples like ListCVAE (Jiang et al., 2018) and PivotCVAE (Liu et al., 2021) use VAEs to generate entire recommendation lists rather than solely ranking individual items.
@@ 2.2.2. Self-Attentive Auto-Regressive Models
Self-attentive models replace the recurrent unit with self-attention and related modules, inspired by transformers (Vaswani et al., 2017). This group of models can be used in session-based recommendation and sequential recommendation (Kang and McAuley, 2018; Wu et al., 2020; Lin et al., 2020; Petrov and Macdonald, 2023), next-basket or bundle prediction (Yu et al., 2020), and model attacking (Yue et al., 2021). Meanwhile, the benefits of self-attentive models are that they handle long-term dependencies better than RNNs and enable parallel training (Vaswani et al., 2017). Additionally, self-attentive models are the de-facto option for pre-trained models (Devlin et al., 2019) and large language models (Brown et al., 2020; Wei et al., 2022; Bubeck et al., 2023), which is gaining traction in RS. More details about using such language models for recommendations will be discussed in Section 3.
@@ 2.4. Diffusion Models
Diffusion models  (Sohl-Dickstein et al., 2015) generate outputs through a two-step process: (1) corrupting inputs into noise via a forward process, and (2) learning to recover the original inputs from the noise iteratively in a reverse process. Their impressive generative capabilities have attracted growing interest from the RS community.
First, a group of works (Wang et al., 2023c; Walker et al., 2022) learns users’ future interaction probabilities through diffusion models. For example, DiffRec (Wang et al., 2023c) predicts users’ future interactions using corrupted noises from the users’ historical interactions. Second, another group of works (Liu et al., 2023d; Wu et al., 2023a) focuses on diffusion models for training sequence augmentation, showing promising results in alleviating the data sparsity and long-tail user problems in sequential recommendation.
@@ 3. Large Language Models in Recommendation
While language has been leveraged by content-based RS for over three decades (Lops et al., 2011), the advent of pretrained LLMs and their emergent abilities for generalized, multi-task natural language (NL) reasoning (Brown et al., 2020; Wei et al., 2022; Bubeck et al., 2023) has ushered in a new stage of language-based recommendation. Critically, NL constitutes a unified, expressive, and interpretable medium that can represent not only item features or user preferences, but also user-system interactions, recommendation task descriptions, and external knowledge (Geng et al., 2022). For instance, items are often associated with rich text including titles, descriptions, semi-structured textual metadata, and reviews. Similarly, user preferences can be articulated in NL in many forms, such as reviews, search queries, liked item descriptions, and dialogue utterances.
Pretrained LLMs provide new ways to exploit this textual data: recent research (e.g., (Sanner et al., 2023; Sileo et al., 2022; Geng et al., 2022; He et al., 2023; Friedman et al., 2023)) has shown that in many domains, LLMs have learned useful reasoning abilities for making and explaining item recommendations based on user preferences as well as facilitating conversational recommendation dialogues. As discussed below, these pretrained abilities can be further augmented through prompting (e.g., (Sanner et al., 2023; Sileo et al., 2022; Liu et al., 2023c)), fine-tuning (e.g., (Geng et al., 2022; Harte et al., 2023; Kang et al., 2023; Zhang et al., 2023b)), retrieval (e.g., (Hou et al., 2023; Wang and Lim, 2023; Dai et al., 2023; Friedman et al., 2023; Kemper et al., 2024)), and other external tools (e.g., (Wang et al., 2022b; Friedman et al., 2023; Zeng et al., 2024).
We next proceed to survey the developments in LLM-based RS’s, first discussing encoder-only LLMs for dense retrieval and cross-encoding (Section 3.1) followed by generative NL recommendation and explanation with sequence-to-sequence (seq2seq) LLMs (Section 3.2). We then review the complementary use of RS and LLMs covering RAG (Section 3.3) and LLM-based feature extraction (Section 3.4), before concluding with a review of conversational recommendation methods (Section 3.5).
@@ 3.1.1. Recommendation as Dense Retrieval
A common task is to retrieve the most relevant items given a NL preference statement using item texts, for which dense retrieval has become a key tool. Dense retrievers (Fan et al., 2022) produce a ranked list of documents given a query by evaluating the similarity (e.g., dot product or cosine similarity) between encoder-only LLM document embeddings and the query embedding. They are highly scalable tools (especially when used with approximate search libraries like FAISS [^1] https://github.com/facebookresearch/faiss) because documents and queries are encoded separately, allowing for dense vector indexing of documents before querying. To use dense retrieval for recommendation (Penha and Hauff, 2020), first, a component of each item’s text content, such as its title, description, reviews, etc., is treated as a document and a dense item index is constructed. Then, a query is formed by some NL user preference description, for instance: an actual search query, the user’s recently liked item titles, or a user utterance in a dialogue.
Several recent works explore recommendation as standard dense retrieval with retrievers that are off-the-shelf (Penha and Hauff, 2020; Harte et al., 2023; Zhang et al., 2023c) and fine-tuned (Mysore et al., 2023; Li et al., 2023c; Hou et al., 2023). More complex dense retrieval methods include review-based retrieval with contrastive BERT fine-tuning (Abdollah Pour et al., 2023) and multi-aspect query decomposition (Korikov et al., 2024), and the use of a second-level encoder to fuse the embedding of a user’s recently liked items into a user embedding before scoring (Wu et al., 2019; Li et al., 2022b).